ruehsjam2004:

Roberto earns $6.50 per hour. He worked 32 hours at the regular rate, 8 hours at a time and a half, and 8 at double time. How much did he earn?

ruehsjam2004:

Any feedback would be greatly appreciated. Thanks!

kacchan:

is it a multiple choice or no?

ruehsjam2004:

No

ruehsjam2004:

I wish

kacchan:

oof

ruehsjam2004:

Tell me about it

kacchan:

give me a minute

zissette:

i know it

kacchan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zissette i know it \(\color{#0cbb34}{\text{End of Quote}}\) if you know it say it then

zissette:

ok give me a sec

zissette:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kacchan \(\color{#0cbb34}{\text{Originally Posted by}}\) @zissette i know it \(\color{#0cbb34}{\text{End of Quote}}\) if you know it say it then \(\color{#0cbb34}{\text{End of Quote}}\) so you do 6.5 times the number of hours

kacchan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zissette \(\color{#0cbb34}{\text{Originally Posted by}}\) @kacchan \(\color{#0cbb34}{\text{Originally Posted by}}\) @zissette i know it \(\color{#0cbb34}{\text{End of Quote}}\) if you know it say it then \(\color{#0cbb34}{\text{End of Quote}}\) so you do 6.5 times the number of hours \(\color{#0cbb34}{\text{End of Quote}}\) don't tell me this i did not ask this question -_-

zissette:

ok

kacchan:

$390 First Roberto works 32 hours at 6.50 per hour. At that time, he earn 32*6.5 = $208. Then he works 8 hours at a time and a half. This means that his pay rate is 50% more, or 6.5*1.5 = $9.75 per hour. Since he worked 8 hours at this pay rate, he earns 8*9.75 = $78. And then he works 8 hours double time. This means his pay rate was doubled during that duration. 6.5*2 = 13, 13*8 = $104. Total: 208 + 78 + 104 = $390

ruehsjam2004:

Thank you so much! You guys are awesome!!!🥰

kacchan:

yw